# Knowledge Editing with Subspace-Aware Key-Value Mappings

### v1

@@ 5 Experiments
Method Counterfact zsRE S Eff. $\uparrow$ Gen. $\uparrow$ Spe. $\uparrow$     Flu. $\uparrow$ Con. $\uparrow$ GC $\uparrow$     S Eff. $\uparrow$ Gen. $\uparrow$ Spe. $\uparrow$ LLama3 Pre-edit 0.0 0.0 0.0 100.0     634.9 20.9 63.4     45.1 35.9 34.8 100.0 FT-L 1.9 9.9 1.4 1.3     438.5 19.2 6.2     43.2 34.8 34.0 88.0 MEND 0.0 0.0 0.0 1.3     519.5 0.6 0.0     0.0 0.0 0.0 89.7 MEMIT 48.3 76.2 74.0 28.2     628.9 36.7 60.8     68.3 89.3 85.7 47.5 PMET 37.6 56.6 56.0 22.6     608.8 33.8 50.1     69.8 89.2 83.9 50.4 AlphaEdit 55.8 97.3 88.7 31.0     633.6 38.6 62.2     73.7 93.5 88.7 53.3 SUIT 86.8 99.7 90.3 74.2     631.2 38.2 63.0     81.6 95.2 85.7 68.5 GPT-J Pre-edit 0.0 0.0 0.0 100.0     621.1 23.9 24.3     35.4 27.2 26.3 100.0 FT-L 13.3 64.9 46.7 5.3     334.2 12.2 24.2     28.4 69.3 60.6 13.4 MEND 0.0 0.0 0.0 0.0     515.0 2.7 0.0     0.6 0.4 0.4 80.0 MEMIT 60.2 92.0 90.4 35.8     617.4 48.5 20.1     88.6 96.8 90.8 79.9 PMET 57.4 84.6 84.5 35.0     618.5 44.8 19.1     85.9 95.1 89.2 75.8 AlphaEdit 73.0 98.3 95.0 49.0     621.8 49.9 19.5     96.9 99.1 92.4 99.4 SUIT 82.3 98.6 93.3 64.1     619.4 49.4 20.4     95.9 99.7 89.5 99.3 Qwen2.5 Pre-edit 0.0 0.0 0.0 100.0     625.5 21.9 28.9     29.5 22.5 21.2 100.0 FT-L 10.3 47.9 31.7 4.2     476.7 4.5 0.0     7.3 18.1 15.6 3.4 MEND 0.0 0.0 0.0 0.0     466.3 0.1 0.0     0.0 0.0 0.0 85.3 MEMIT 22.6 83.0 84.0 9.2     622.1 37.0 34.8     72.7 82.9 72.2 65.1 PMET 32.6 67.5 65.9 16.1     545.2 27.5 14.8     60.2 71.9 65.5 48.4 AlphaEdit 67.8 97.1 91.6 43.4     626.2 41.0 28.1     89.6 97.5 86.3 85.8 SUIT 85.7 99.5 86.8 74.4     626.2 37.4 30.8     88.2 93.9 76.9 96.6
@@ 5.1 Experimental Setup
Our objective is editing, the task of changing $(s,r,o)$ to $(s,r,o^{\ast})$. Since editing presupposes the model’s prior knowledge of a fact, we exclusively evaluate on instances where the model’s argmax prediction is $o$ for the original prompt $(s,r)$ as well as its corresponding paraphrase and neighborhood prompts. This condition is not applied to zsRE, which contrasts with declarative datasets like CounterFact. Its question-based prompts lead the model to predict sentence-starters (e.g., “The”) rather than the immediate object, making the argmax check unsuitable. In Table 1, we present the results of applying 1000 edits sequentially in 10 batches of 100. For this evaluation, an edit is considered successful based on a generation-based criterion, which counts an edit as successful only if $o^{\ast}$ is the argmax prediction. In the Section F.2, we also report results with the less demanding probability-based criterion ($P(o^{\ast})>P(o)$), common in prior work.
@@ 6 Analysis
To further investigate SUIT, we conducted a series of analyses. We demonstrate its effectiveness in reducing perturbation at the subject’s last-token. We then empirically validate our hypotheses regarding the subspaces identified for computing the key and residual vectors. Finally, we present a hyperparameter tradeoff analysis and an ablation study. All analyses were performed using the Llama-3-Instruct model. For experiments requiring edited models, we used models edited with 10 batches of 100 edits each from the CounterFact.
@@ 6.1 Reducing Entity’s Last Token Perturbation
Figure 3 shows the $L_{2}$ norm of difference in the residual streams of the final edited layer between the original model and models edited using MEMIT, AlphaEdit, and SUIT. This difference, representing the perturbation at each token, is visualized through color intensity on a sample paragraph from the Wikinews Article Dataset. The figure demonstrates that, compared to other tokens, all methods generally exhibit a more notable perturbation at the last token of the subject entity (e.g., “OS” in “NOS”, “ner” in “Donner”). Notably, among these methods, the perturbation from SUIT is visibly less pronounced than that of MEMIT and AlphaEdit. Further examples confirming this pattern are provided in the Appendix H.
@@
Figure 3: Norm differences of MLP outputs at the last token position across methods.
@@ 6.2.2 Analysis for $\mathbf{w}_{1}$, $\mathbf{w}_{2}$
We further analyzed the individual roles of $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ to test if $\mathbf{w}_{1}$ primarily increases the logit of the new object $o^{\ast}$ while $\mathbf{w}_{2}$ primarily decreases the logit of the old object $o$. To do this, we decomposed our update vector $\delta^{\prime}$ into its components along these directions, $\Delta\mathbf{w}_{1}=(\mathbf{h}^{\top}\mathbf{w}_{2}-\mathbf{h}^{\top}\mathbf{w}_{1})\mathbf{w}_{1}$, $\Delta\mathbf{w}_{2}=(\mathbf{h}^{\top}\mathbf{w}_{1}-\mathbf{h}^{\top}\mathbf{w}_{2})\mathbf{w}_{2}$. We then incrementally added each component to the residual stream $\mathbf{h}$ by scaling it with an interpolation factor $k\in[0,1]$ and observed the logits for both objects.
@@ 6.3 Tradeoff and Ablation Study
Regarding the energy threshold $\tau_{\text{energy}}$ , the efficacy score tended to remain stable as the threshold increased. The generation score showed a sweet spot around 0.3-0.4, which appears to be effective for identifying a suitable subspace $\mathcal{K}_{s}^{\perp}$. As $\tau_{\text{energy}}$ increased further, the subspace considered for editing tended to narrow. This created a clear tradeoff: the generation score decreased as the model became less fitted to the edits, while the specificity score increased.
@@
Method Eff. Gen. Spe.
SUIT ($k$ + $\delta$) 99.7 90.3 74.2
$k$-only 96.4 77.9 74.7
$\delta$-only 99.7 83.8 44.6
To isolate the contribution of each of our proposed components, we conducted an ablation study on the key vector computation ($\mathbf{k}$-only) and the residual vector computation ($\mathbf{\delta}$-only), with the results summarized in Table 5. The study reveals a powerful synergy between the two components. Our full method not only preserves the high efficacy of the $\mathbf{\delta}$-only method and the strong specificity of the $\mathbf{k}$-only method but also achieves a generalization score that significantly surpasses both. This result validates the effectiveness of our integrated approach.
@@ Appendix B The Regularization Term $\mathcal{R}$
The second term is a weight decay penalty on the L2 norm of the residual vector $\tilde{\delta}$. This term encourages the model to find a smaller solution for $\tilde{\delta}$. The hyperparameter $\lambda_{\text{KL}}$ controls the strength of this penalty.
@@ D.2 Hyperparameter Settings
We adopt a unified configuration across locate–then–edit methods (ROME, MEMIT, PMET, RECT, PRUNE, NSE) as well as AlphaEdit and SUIT, and only deviate where method-specific constraints apply. Unless noted otherwise, we set v_weight_decay= 0.5, and kl_factor= 0.0625. For model-specific layer selections, we use indices {4, 5, 6, 7, 8} for Llama3-8B-Insctruct, {3, 4, 5, 6, 7, 8} for Gpt-J-6b, and {4, 5, 6, 7, 8} for Qwen2.5-7B-Instruct; an exception is ROME, which always edits a single target layer (index 5). SUIT does not use v_weight_decay or kl_factor. In addition, both AlphaEdit and SUIT employ nullspace_threshold= $2\times 10^{-2}$ and L2= 10. For SUIT, we fix the method-internal hyperparameters $\tau_{\text{energy}}=0.4$ and $\lambda=0.3$ for all models. MEND and FT-L/FT-W are tuned following their original papers. Hyperparameter choices follow those specified in the original papers; when not provided, we default to the EasyEdit open-source settings Wang et al. (2024a).
@@ E.1 CounterFact Metrics
Here, $o$ denotes the first token generated by the pre-edit model for $O(x_{i})$ rather than the dataset’s gold answer; we adopt this choice because the language model often does not reproduce the zsRE-provided $o$. Average token-level accuracy on neighborhood prompts:
@@ F.1 Extended Baseline Comparisons
Method Counterfact Eff. $\uparrow$ Gen. $\uparrow$ Spe. $\uparrow$ Flu. $\uparrow$     Con. $\uparrow$ GC $\uparrow$ LLaMA3 Pre-edit 0.0 0.0 100.0 634.9     20.9 62.3 FT-W 4.0 2.9 43.5 634.4     21.4 60.5 ROME 0.0 0.2 0.0 481.5     4.2 0.0 RECT 81.6 72.4 36.1 634.4     35.3 60.4 PRUNE 43.3 39.3 17.7 590.4     33.9 45.0 NSE 1.4 5.8 62.2 609.6     23.1 60.9 SUIT 99.7 90.3 74.2 631.2     38.2 61.8 GPT - J Pre-edit 0.0 0.0 100.0 621.1     23.9 18.6 FT-W 12.3 2.4 49.0 613.4     25.7 36.1 ROME 0.1 0.2 0.2 407.2     4.2 0.0 RECT 92.9 85.8 44.4 625.0     47.6 14.9 PRUNE 51.8 56.0 16.9 504.6     29.4 29.4 NSE 0.8 12.6 54.1 608.0     34.2 27.5 SUIT 98.6 93.3 64.1 619.4     49.4 17.8 Qwen2.5 Pre-edit 0.0 0.0 100.0 625.5     21.9 20.8 FT-W 47.9 31.7 4.2 476.7     4.5 0.0 ROME 51.6 33.7 14.2 440.1     15.6 0.6 RECT 86.3 85.9 42.3 625.8     37.7 59.9 PRUNE 28.2 30.7 7.7 588.1     30.4 6.0 NSE 0.0 0.0 99.5 625.6     21.7 39.3 SUIT 99.5 86.8 74.4 626.2     37.4 23.7

### v2

@@ 5 Experiments
Method Counterfact zsRE S $\uparrow$ Eff. $\uparrow$ Gen. $\uparrow$ Spe. $\uparrow$     Flu. $\uparrow$ Con. $\uparrow$ GC $\uparrow$     S $\uparrow$ Eff. $\uparrow$ Gen. $\uparrow$ Spe. $\uparrow$ LLama3 Pre-edit 0.0 0.0 0.0 100.0     634.9 20.9 63.4     45.1 35.9 34.8 100.0 FT-L 1.9 9.9 1.4 1.3     438.5 19.2 6.2     43.2 34.8 34.0 88.0 MEND 0.0 0.0 0.0 1.3     519.5 0.6 0.0     0.0 0.0 0.0 89.7 MEMIT 48.3 76.2 74.0 28.2     628.9 36.7 60.8     68.3 89.3 85.7 47.5 PMET 37.6 56.6 56.0 22.6     608.8 33.8 50.1     69.8 89.2 83.9 50.4 AlphaEdit 55.8 97.3 88.7 31.0     633.6 38.6 62.2     73.7 93.5 88.7 53.3 SUIT 86.8 99.7 90.3 74.2     631.2 38.2 63.0     81.6 95.2 85.7 68.5 GPT-J Pre-edit 0.0 0.0 0.0 100.0     621.1 23.9 24.3     35.4 27.2 26.3 100.0 FT-L 13.3 64.9 46.7 5.3     334.2 12.2 24.2     28.4 69.3 60.6 13.4 MEND 0.0 0.0 0.0 0.0     515.0 2.7 0.0     0.6 0.4 0.4 80.0 MEMIT 60.2 92.0 90.4 35.8     617.4 48.5 20.1     88.6 96.8 90.8 79.9 PMET 57.4 84.6 84.5 35.0     618.5 44.8 19.1     85.9 95.1 89.2 75.8 AlphaEdit 73.0 98.3 95.0 49.0     621.8 49.9 19.5     96.9 99.1 92.4 99.4 SUIT 82.3 98.6 93.3 64.1     619.4 49.4 20.4     95.9 99.7 89.5 99.3 Qwen2.5 Pre-edit 0.0 0.0 0.0 100.0     625.5 21.9 28.9     43.7 34.4 33.8 100.0 FT-L 10.3 47.9 31.7 4.2     476.7 4.5 0.0     17.0 12.5 11.7 89.7 MEND 0.0 0.0 0.0 0.0     466.3 0.1 0.0     0.0 0.0 0.0 10.6 MEMIT 22.6 83.0 84.0 9.2     622.1 37.0 34.8     76.2 96.2 91.4 55.4 PMET 32.6 67.5 65.9 16.1     545.2 27.5 14.8     58.3 82.1 76.3 38.2 AlphaEdit 67.8 97.1 91.6 43.4     626.2 41.0 28.1     73.9 94.5 89.8 52.9 SUIT 85.7 99.5 86.8 74.4     626.2 37.4 30.8     89.3 97.1 90.5 81.7
@@ 5.1 Experimental Setup
Our objective is editing, the task of changing $(s,r,o)$ to $(s,r,o^{\ast})$. Since editing presupposes the model’s prior knowledge of a fact, we exclusively evaluate on instances where the model’s argmax prediction is $o$ for the original prompt $(s,r)$ as well as its corresponding paraphrase and neighborhood prompts. This condition is not applied to zsRE, which contrasts with declarative datasets like CounterFact. Its question-based prompts lead the model to predict sentence-starters (e.g., “The”) rather than the immediate object, making the argmax check unsuitable (details in Section D.3). In Table 1, we present the results of applying 1000 edits sequentially in 10 batches of 100. For this evaluation, an edit is considered successful based on a generation-based criterion, which counts an edit as successful only if $o^{\ast}$ is the argmax prediction. In the Section F.2, we also report results with the less demanding probability-based criterion ($P(o^{\ast})>P(o)$), common in prior work.
@@ 6 Analysis
To further investigate SUIT, we conducted a series of analyses. We demonstrate its effectiveness in reducing perturbation at the subject’s last-token. We then empirically validate our hypotheses regarding the subspaces identified for computing the key and residual vectors. Finally, we present a hyperparameter tradeoff analysis and an ablation study. All analyses were performed using the LLaMA3-Instruct (8B) model. For experiments requiring edited models, we used models edited with 10 batches of 100 edits each from the CounterFact.
@@ 6.1 Reducing Entity’s Last Token Perturbation
Figure 3 shows the $L_{2}$ norm of difference in the residual streams of the final edited layer between the original model and models edited using MEMIT, AlphaEdit, and SUIT. This difference, representing the perturbation at each token, is visualized through color intensity on a sample paragraph from the Wikinews Article Dataset. The figure demonstrates that, compared to other tokens, all methods generally exhibit a more notable perturbation at the last token of the subject entity (e.g., “OS” in “NOS”, “ner” in “Donner”). Notably, among these methods, the perturbation from SUIT is visibly less pronounced than that of MEMIT and AlphaEdit. Further examples confirming this pattern are provided in the Appendix K.
@@
Figure 3: Norm differences of MLP outputs at the subject entity’s last token position across methods.
@@ 6.2.2 Analysis for $\mathbf{w}_{1}$, $\mathbf{w}_{2}$
We further analyzed the individual roles of $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ to test if $\mathbf{w}_{1}$ primarily increases the logit of the new object $o^{\ast}$ while $\mathbf{w}_{2}$ primarily decreases the logit of the old object $o$. To do this, we decomposed our update vector $\delta^{\prime}$ into its components along these directions, $\Delta\mathbf{w}_{1}=(\mathbf{h}^{\top}\mathbf{w}_{2}-\mathbf{h}^{\top}\mathbf{w}_{1})\mathbf{w}_{1}$, $\Delta\mathbf{w}_{2}=(\mathbf{h}^{\top}\mathbf{w}_{1}-\mathbf{h}^{\top}\mathbf{w}_{2})\mathbf{w}_{2}$.
@@
We then incrementally added each component to the residual stream $\mathbf{h}$ by scaling it with an interpolation factor $k\in[0,1]$ and observed the logits for both objects.
@@ 6.3 Tradeoff and Ablation Study
Regarding the energy threshold $\tau_{\text{energy}}$, the Efficacy score remains largely stable as the threshold increases. The trade-off between Generalization and Specificity is most mitigated around $\tau_{\text{energy}}\approx 0.3$$0.4$, suggesting that this moderate range is effective for identifying a suitable editing subspace $\mathcal{K}s^{\perp}$. Accordingly, we select $\tau_{\text{energy}}$ by maximizing the harmonic-mean score $S$. To examine whether this selection range generalizes, we additionally analyzed cumulative energy curves across multiple models and observed highly consistent patterns (details in Appendix H). This consistency indicates that the selected $\tau_{\text{energy}}$ range remains stable when extended to other models.
@@
Method S Eff. Gen. Spe.
SUIT ($k$ + $\delta$) 86.8 99.7 90.3 74.2
$k$-only 82.2 96.4 77.9 74.7
$\delta$-only 68.3 99.7 83.8 44.6
Our methodology comprises two components: key vector computation ($\mathbf{k}$-only) and the residual vector computation ($\mathbf{\delta}$-only). To verify that SUIT’s performance stems from the synergy between these two elements rather than a single factor, we conducted an ablation study, with results summarized in Table 5. Compared to the AlphaEdit results on LLaMA in Table 1, both k-only and $\delta$-only achieve higher S scores individually, and applying both together yields an even greater improvement. This demonstrates a clear synergy between the two components. When combined, they substantially enhance both efficacy and specificity, confirming the effectiveness of the integrated approach.
@@ Appendix B The Regularization Term $\mathcal{R}$
The second term is a weight decay penalty on the L2 norm of the residual vector $\tilde{\delta}$. This term encourages the model to find a smaller solution for $\tilde{\delta}$. The hyperparameter $\lambda_{\text{WD}}$ controls the strength of this penalty.
@@ D.2 Hyperparameter Settings
We adopt a unified configuration across locate–then–edit methods (ROME, MEMIT, PMET, RECT, PRUNE, NSE) as well as AlphaEdit and SUIT, and only deviate where method-specific constraints apply. Unless noted otherwise, we set v_weight_decay= 0.5, and kl_factor= 0.0625. For model-specific layer selections, we use indices {4, 5, 6, 7, 8} for LLaMA3-Instruct (8B), {3, 4, 5, 6, 7, 8} for GPT-J (6B), and {4, 5, 6, 7, 8} for Qwen2.5-Instruct (7B); an exception is ROME, which always edits a single target layer (index 5). SUIT does not use v_weight_decay or kl_factor. In addition, both AlphaEdit and SUIT employ nullspace_threshold= $2\times 10^{-2}$ and L2= 10. For SUIT, we used $\tau_{\text{energy}}=0.4$ and $\lambda=0.3$ as hyperparameters. As shown in Figure 5, the overall performance does not vary substantially when $\tau_{\text{energy}}$ is set within the range of 0.4–0.5. Likewise, changing $\lambda$ between 0.3 and 0.4 leads to no noticeable difference in overall performance. MEND and FT-L/FT-W are tuned following their original papers. Hyperparameter choices follow those specified in the original works; when not provided, we default to the EasyEdit open-source settings Wang et al. (2024a).
D.3 zsRE Dataset Details
Unlike declarative datasets such as CounterFact, zsRE uses question-style prompts, which makes an argmax-based pre-knowledge check unreliable. In particular, answers to zsRE questions frequently begin with function words or pronouns rather than the target object span. As a result, the model’s first-token argmax is often unrelated to the gold object (e.g., common starters include “the” 4,902 times, “a” 445 times, “he” 388 times, “she” 89 times). This violates the core assumption behind the argmax criterion—that the argmax at the object position reflects whether the model already knows $(s,r,o)$.
Moreover, the model often responds with a full sentence instead of a short object-only span, even when the underlying fact is known. For example, given a prompt like “What airport is China United Airlines part of?”, the gold object is “Beijing Nanyuan Airport,” but the model may produce a longer answer such as “China United Airlines is part of Beijing Capital International Airport.” In such cases, a strict argmax pre-knowledge condition would incorrectly reject the instance or mark it as misaligned, despite the model engaging with the correct relation.
For these reasons, applying the same pre-knowledge restriction as in CounterFact would discard a large fraction of valid zsRE instances and distort evaluation. We therefore evaluate zsRE without this condition.
@@ E.1 CounterFact Metrics
Here, $o$ denotes the first token generated by the model immediately before performing the current edit batch for $O(x_{i})$, which we use as the preservation reference, rather than the dataset’s gold answer; we adopt this choice because the language model often does not reproduce the zsRE-provided $o$.
In fact, as discussed in Appendix E.2.1, gold-referenced partial-credit scores are already very low pre-edit on zsRE, making them a weak signal for preservation. Average token-level accuracy on neighborhood prompts:
@@
E.2.1 Limitations of Standard Specificity Measurement on zsRE
Prior work evaluating zsRE (Fang et al., 2025; Meng et al., 2023a; b) measures Specificity via token-level partial credit to the dataset’s ground-truth answer. However, on zsRE this standard choice is often uninformative: even before editing, the model already receives very low partial-credit scores, so small post-edit differences are difficult to interpret as genuine preservation failures. The following results are obtained under the same sequential editing setup as Table 1.
Method LLaMA GPT-J
Pre-edit 32.0 27.0
AlphaEdit 32.7 28.2
SUIT 32.7 27.8
Table 6: zsRE Specificity measured with token-level partial credit to the dataset ground-truth answers in the sequential setting (1,000 samples; 10 batches of 100 edits).
To make zsRE’s Specificity more informative, we instead use the model’s own generated tokens as the reference to preserve. Concretely, in the sequential setting, we record the model’s generated token sequence for each batch’s specificity prompts immediately before applying that batch edit, and measure how much these answers change due to the batch. Table 1 is computed under this setup.
E.2.2 Extended Evaluation of zsRE Specificity under a Stricter Test
To better demonstrate SUIT’s strength on zsRE Specificity, we run an additional, stricter evaluation. The batch-level setting reported in Table 1 is relatively lenient because it uses the model’s answer immediately before each batch as the preservation reference, capturing only the incremental drift introduced by that batch edit. Here, we instead fix the model’s original pre-edit answer as an reference, and after completing all sequential edits, check whether its first 1 to 4 tokens are exactly preserved. Under this stricter test, SUIT preserves prior answers more reliably than AlphaEdit.
Method Specificity tok1 tok2 tok3 tok4 LLaMA Pre-edit 100.0 100.0 100.0 100.0 AlphaEdit 31.2 22.1 19.8 18.3 SUIT 51.9 43.2 41.2 39.8 GPT - J Pre-edit 100.0 100.0 100.0 100.0 AlphaEdit 94.1 93.8 68.8 65.3 SUIT 99.5 99.3 83.7 80.3
Table 7: Extended zsRE Specificity evaluation under a stricter preservation test, using the same sequential editing setup as Table 1. tokkk denotes exact preservation of the first $k$ tokens of the original pre-edit answer after all edits.
@@ F.1 Extended Baseline Comparisons
Table 8 presents results for additional extended baselines that were not reported in Table 1.
Method Counterfact Eff. $\uparrow$ Gen. $\uparrow$ Spe. $\uparrow$ Flu. $\uparrow$     Con. $\uparrow$ GC $\uparrow$ LLaMA Pre-edit 0.0 0.0 100.0 634.9     20.9 62.3 FT-W 4.0 2.9 43.5 634.4     21.4 60.5 ROME 0.0 0.2 0.0 481.5     4.2 0.0 RECT 81.6 72.4 36.1 634.4     35.3 60.4 PRUNE 43.3 39.3 17.7 590.4     33.9 45.0 NSE 1.4 5.8 62.2 609.6     23.1 60.9 SUIT 99.7 90.3 74.2 631.2     38.2 61.8 GPT - J Pre-edit 0.0 0.0 100.0 621.1     23.9 18.6 FT-W 12.3 2.4 49.0 613.4     25.7 36.1 ROME 0.1 0.2 0.2 407.2     4.2 0.0 RECT 92.9 85.8 44.4 625.0     47.6 14.9 PRUNE 51.8 56.0 16.9 504.6     29.4 29.4 NSE 0.8 12.6 54.1 608.0     34.2 27.5 SUIT 98.6 93.3 64.1 619.4     49.4 17.8 Qwen2.5 Pre-edit 0.0 0.0 100.0 625.5     21.9 20.8 FT-W 47.9 31.7 4.2 476.7     4.5 0.0 ROME 51.6 33.7 14.2 440.1     15.6 0.6 RECT 86.3 85.9 42.3 625.8     37.7 59.9 PRUNE 28.2 30.7 7.7 588.1     30.4 6.0 NSE 0.0 0.0 99.5 625.6     21.7 39.3 SUIT 99.5 86.8 74.4 626.2     37.4 23.7